# Photoemission tomography of excitons in 2D systems: momentum-space signatures of correlated electron-hole wave functions

### v1

@@ II.2 Pump-induced effects on exciton photoemission
In the preceding derivation, we assumed the system was in an eigenstate of the BSE Hamiltonian, $|m,\boldsymbol{Q}\rangle$, whose real-space representation is shown in Equation 1. The resulting expression for the photoelectron intensity (Equation II.1) applies to such an eigenstate and does not consider how the system was initially prepared by the pump pulse. In general however, the pump pulse typically generates a superimposed state—a linear combination of eigenstates—which necessitates an adaptation of our formalism, as detailed below. Here, we restrict the discussion to the optical limit ($\boldsymbol{Q}\rightarrow 0$), allowing us to set $|m,\boldsymbol{Q}=0\rangle=:|m\rangle$.
The system’s excitation probability due to the pump pulse—characterized by its polarization vector $\boldsymbol{E}$ and photon energy $\Omega$—is related to the optical spectrum, which is commonly described by the imaginary part of the macroscopic dielectric function [2]:
$\Im\{\epsilon\}(\Omega)=\frac{16\pi\mathrm{e}^{2}}{\Omega^{2}}\sum_{m}\left|\langle 0|\boldsymbol{E}\cdot\hat{\boldsymbol{v}}|m\rangle\right|^{2}\delta(\Omega-\Omega^{(m)}).$ (17)
Here $|0\rangle$ denotes the ground state, $\hat{\boldsymbol{v}}=\mathrm{i}[\hat{H},\hat{\boldsymbol{r}}$] the single-particle velocity operator and we assume $|\boldsymbol{E}|=1$. In the same spirit, we can sum excitons in the energy-degenerate subspace of the BSE Hamiltonian, in order to prepare an initial superimposed state, as created by the pump field. For a small enough energy difference $\eta>0$, which spans the degenerate subspace around excitation energy $\Omega_{l}$, we write the superimposed state $|S_{l}\rangle$ as a linear combination of eigenstates,
$|S_{l}\rangle=\sum_{m}c_{m}^{l}|m\rangle,$ (18)
with expansion coefficients $c_{m}^{l}$ depending on the polarization of the pump via the optical transition matrix elements:
$c_{m}^{l}=\langle 0|\boldsymbol{E}\cdot\hat{\boldsymbol{v}}|m\rangle\;\delta\left(|\Omega^{(m)}-\Omega_{l}|-\eta\right).$ (19)
As a consequence, for photoemission intensity from a superimposed state $|S_{l}\rangle$ also the transition density matrix elements are changed and thus the final results in Equation II.1 must be modified accordingly (for $\boldsymbol{Q}=0$):
@@
$\cdot\delta\left(\omega+\Omega_{l}-\varepsilon_{v,\boldsymbol{q}}-E_{\mathrm{kin}}\right).$ (20)
In 2D systems excitons typically occur as degenerate pairs, and only their superposition should be considered as the physical exciton, which then, for instance, also respects the symmetries of the system [40, 41]. In contrast to those eigenstates, the pump pulse in general breaks the system’s symmetry and may create superimposed exciton states that depend on the pump polarization. As we will show in Paragraph II.2, such pump polarization effects have a considerable imprint on exciton photoemission. It should be noted, however, that our formalism is only suitable for degenerate energies (the condition of $\eta$ being small) and not for the case of a detuned pump pulse, i.e. excitons excited off resonance, and that the above discussion required $\boldsymbol{Q}=0$.
@@ III.1 Calculation of hBN quasiparticle and optical properties
In the following, we exemplify our formalism by the simulation of excitonic photoemission in monolayer hBN as a prototypical 2D system. As a starting point, we obtain wave functions and energy levels from DFT calculations with Quantum Espresso [42, 43, 44], using the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [45] and a lattice constant of 2.508 $\mathrm{\AA }$, see Figure 1 (a). To compensate for the systematic underestimation of the band gap within semi-local DFT, we apply the $G_{0}W_{0}$ method to correct the energy levels. Utilizing the BerkeleyGW code [1, 2, 46], we obtain a direct (indirect) band gap of 7.6 eV (7.4 eV), which is comparable to literature values from similar methods, see e.g. Reference [47] for a comparison. We show the $G_{0}W_{0}$ band structure in Figure 1 (b) and detail the computational methods in Appendix B.
Figure 2: Optical excitation of monolayer hBN. (a) Independent-particle spectrum, based on a $G_{0}W_{0}$ calculation in dark blue and the BSE spectrum in pink. In the latter, $S_{1}$ and $S_{2}$ energies are marked by arrows. Panels (b) and (c) show the $G_{0}W_{0}$ band structure with overlays based on the BSE eigenvector composition of the $S_{1}$ (blue) and $S_{2}$ (green) excitations of the system, respectively.
Subsequently, we use the DFT wave functions and $G_{0}W_{0}$ energy levels, and solve the BSE within the TDA and for $\boldsymbol{Q}=0$. This yields the excitation energies $\Omega_{m}$ and the BSE eigenvectors $|m,\boldsymbol{Q}=0\rangle\equiv|m\rangle$, which we use to simulate optical spectra via Equation 17. As shown in Figure 2 (a), the BSE spectrum features a range of excitonic peaks below the band gap. These are often interpreted as hydrogen-like solutions of the Wannier equation [48], though they have also been found to deviate from that model [40]. For comparison, we include the optical spectrum computed within the independent-particle approximation (IPA), shown as a dashed blue curve in Figure 2 (a). In the following, we focus on the two lowest excitons, located at 5.3 eV ($S_{1}$) and 6.2 eV ($S_{2}$).
Both the $S_{1}$ and $S_{2}$ superimposed excitons are composed of two degenerate eigenstates of the BSE Hamiltonian. This results in an additional dependence on the pump polarization, as discussed in Paragraph II.2, and which will become relevant at a later stage in Paragraph III.3. For $S_{1}$ at 5.3 eV, the two degenerate valence-conduction transitions occur at $K$ and $K^{\prime}$, respectively, with their momentum-dependent transition weights indicated in Figure 2 (b). The situation is similar for $S_{2}$ at 6.2 eV, albeit with intensities being more spread out along the Brillouin-zone boundary, as can be seen in Figure 2 (c).
@@ III.2 Exciton photoemission in hBN
With the quasiparticle and optical properties at hand, we now illustrate consequences of our theory for exciton photoemission from monolayer hBN. In the remainder of this manuscript, we use a 26.5 eV probe laser with a Gaussian energy broadening of 0.1 eV. For a better insight into the pump-induced effects on the momentum maps, we ignore the trivial effect of probe polarization, stemming from the $|\boldsymbol{A}\boldsymbol{k}|^{2}$ prefactor in Equation II.1. Unlike other theoretical formulations, we only consider excitonic photoemission. As a result, the populated electronic structure and its depopulation due to exciton creation are not accounted for. This is not a significant drawback, since in trARPES experiments the unoccupied part is typically recorded with much lower signal statistics.
Figure 3: ARPES bandmap for the $S_{1}$ exciton in $k_{x}$-direction (cut of the BZ along $\Gamma-K^{\prime}-M_{2}$, see inset with three neighboring BZs). The $G_{0}W_{0}$ bandstructure is overlaid in full blue lines, with the valence band, shifted by $\Omega$, indicated in blue dashed lines.
In Figure 3 we show the simulated bandmap for photoemission from the $S_{1}$ exciton together with the $G_{0}W_{0}$ bandstructure (full blue lines) along the $\Gamma-K^{\prime}-M_{2}$ path (corresponding to the $k_{x}$-direction, see inset). According to Equation II.1, the energy dispersion of exciton photoemission is determined by the hole component of the exciton. Consequently, the exciton’s trARPES signature appears at the photoelectron kinetic energy corresponding to the valence band energy plus the pump energy $\Omega$ (blue dashed line). This finding, which is well-established in the literature [49, 19], enables the estimation of the exciton’s spatial extent from photoemission data under the assumption of an $s$-like wave function [4, 5, 50].
@@
For this purpose, it is expedient to compare momentum maps of exciton photoemission with those for photoemission from the ground state, which we obtain by the standard way of simulating POT [51]. In Figure 4 (a), we present such a map for a kinetic energy slightly below the valence band maximum (VBM) ($E-E_{\mathrm{VBM}}=-0.2$ eV), which is comprised of six triangular features, each around the $KK^{\prime}$ points. This pattern is analogous to the case of graphene, where the triangular features are also broken in the outgoing direction of the Brillouin zone, leading to croissant- or horseshoe-like shapes. This “dark corridor” is due to matrix element effects in the photoemission process and has already been described in great detail for the case of graphene [52, 53, 14]. When assuming a static, charge-transfer induced population of the conduction band, the latter would also be accessible to photoemission from the ground state, which we simulate in panel (b) of Figure 4, using a kinetic energy of 26.6 eV, which is above the conduction band minimum (CBM) ($E-E_{\mathrm{CBM}}=+0.6$ eV). In contrast to the valence band case depicted in panel (a), the features around the Dirac point are now elongated in the directions going to the $MM^{\prime}$ points, which is reasonable when looking at the bandstructure (see e.g. Figure 3). Remarkably, the dark corridor from the metrix element effect is now pointing to the inside of the BZ. Returning now to the simulation of exciton photoemission, we show a momentum map of the $S_{1}$ exciton in Figure 4 (c), using a kinetic energy of 23.9 eV and a pump polarization as indicated by a black arrow in Figure 5. As already discussed in the preceeding paragraph, the exciton inherits the energy dispersion of the hole (here: the highest valence band), but, in light of Equation II.1, follows the momentum distribution given by the Fourier-transformed electron part of the exciton wave function (here: the lowest conduction band). As a direct imprint of the latter, the dark corridor is also pointing to the inside of the BZ, which we can see clearly in Figure 4 (c). Having established the unoccupied part of the electronic structure as another direct marker of exciton photoemission, we will now examine the different contributions to the momentum maps in more detail.
@@
Figure 5: Decomposition of exciton ARPES maps for $S_{1}$ in row (a) and $S_{2}$ in row (b), with black arrows indicating the polarization of the pump laser. First column: Fourier-transformed conduction-band wave functions. Second column: energy-conservation delta-function. Third column: BSE eigenvectors. Fourth column: resulting exciton ARPES maps as procuts of first to third columns. Each map is individually normalized to one. For $S_{1}$ in row (a), the horseshoes point away from $\Gamma$, with minor modulations along their contours. For $S_{2}$ in row (b), four of the six horseshoes are rotated, when compared to the first exciton. This is a direct result of the differences in the momentum-space distribution of their BSE eigenvectors.
@@
Using Equation II.1, we dissect the photoemission maps for both excitons, $S_{1}$ and $S_{2}$, in order to understand the influence of each contributing factor. Since—in this case—both excitons only involve transitions from the highest valence to the lowest conduction band, the sums over $v$ and $c$ in Equation II.1 collapse. Consequently, we are left with a single product of $|\mathcal{F}[\xi]|^{2}$, $\delta_{E}$ and $|X|^{2}$, involving only wave functions from the lowest conduction band, energies of the highest valence band and BSE transition coefficients from the highest valence band to the lowest conduction band. Figure 5 shows these three factors, as well as the full photoemission maps, for the two excitons in row (a) ($S_{1}$) and row (b) ($S_{2}$), respectively. The direction of the pump polarization is indicated in the rightmost columns. For both excitons $S_{1}$ and $S_{2}$, the triangular patterns are generated by the $\delta_{E}$ factor, with $|\mathcal{F}[\xi]|^{2}$ only leading to slight modifications. The maps in the first two columns, $|\mathcal{F}[\xi]|^{2}$ and $\delta_{E}$, are very similar between the two different excitons $S_{1}$ and $S_{2}$, with minor differences arising due to the different photoelectron kinetic energies (23.9 eV for $S_{1}$ and 24.8 eV for $S_{2}$). Therefore, the additional multiplication with the BSE eigenvector can at most lead to intensity changes within the area of the horseshoes. As can be seen from the third column, the BSE eigenvectors of the two excitons differ, leading to the differences in their respective photoemission maps (rightmost column of Figure 5). In the following, those differences are inspected more closely.
@@
For $S_{1}$ in Figure 5 (a), the BSE eigenvector (third column) mostly exhibits a six-fold symmetry. It is centered around the $K$ and $K^{\prime}$ points and varies weakly along the contours of the horseshoes, thus leading only to small intensity variations along them. Contrary, in the case of $S_{2}$ in row (b), the six-fold symmetry is broken by the BSE eigenvector, which exhibits a nodal plane perpendicular to the pump polarization, while being non-zero along the remaining four sides of the hexagonal BZ. Since the pump polarization was chosen along a high-symmetry direction (indicated by black arrows in the rightmost column), it acts as a symmetry axis for the BSE eigenvector and, consequently, for the resulting photoemission map. In contrast to the case of $S_{1}$, the regions around $K$ and $K^{\prime}$ show much more details for $S_{2}$, particularly featuring nodal lines, which, upon multiplication of the BSE eigenvector with the horseshoe pattern produced by $|\mathcal{F}[\xi]|^{2}$ and $\delta_{E}$, suppress photoemission intensity. Unlike $S_{1}$, the regions around $K$ and $K^{\prime}$ for $S_{2}$ exhibit more detailed features, including nodal lines, for which photoemission intensity is suppressed in the product of $X$ with the horseshoe pattern produced by $|\mathcal{F}[\xi]|^{2}$ and $\delta_{E}$. Moreover, the local maxima are shifted off $K$ and $K^{\prime}$. As can be seen in the last column of Figure 5, the combination of all three factors effectively leads to a redistribution of spectral intensity within four out of the six horseshoes. By selectively accessing the different factors in Equation II.1, we are thus able to deconvolute different contributions to the momentum maps. In the general case of multiple transitions to an exciton, however, the sums over different valence and conduction bands remain and, as a result, the above analysis would become less straightforward, especially due to the coherent summation of the electron part. Since in the present case of the hBN model we found the most asymmetry arising due to the BSE eigenvectors in conjunction with the pump polarization, we will discuss this effect in more detail in the following.

### v2

@@ II.2 Pump-induced effects on exciton photoemission
In the preceding derivation, we assumed the system to be in an eigenstate of the BSE Hamiltonian, $|m,\boldsymbol{Q}\rangle$, whose real-space representation is given in Eq. 1. The resulting expression for the photoelectron intensity (Eq. II.1) applies to such an eigenstate and does not consider how the system was initially prepared by the pump pulse. In general however, the pump pulse typically generates a superposition state—a linear combination of the calculated eigenstates—which necessitates an adaptation of our formalism, as detailed below. In preparation for Sec. III, we restrict the discussion to optical excitations ($\boldsymbol{Q}\rightarrow 0$), allowing us to denote the eigenstates of the BSE Hamiltonian as $|m,\boldsymbol{Q}=0\rangle=:|m\rangle$. The superposition state $|S_{l}\rangle$, created by the pump laser, then becomes
$|S_{l}\rangle=\sum_{m}c_{m}^{l}|m\rangle.$ (17)
As mentioned above, the presented formalism Eq. II.1 is only suitable for photoemission from an eigenstate of the BSE Hamiltonian. Assuming that the superposition state $|S_{l}\rangle$, created by the pump field, is a combination of energy-degenerate eigenstates (obtained from solving the BSE), it is itself an eigenstate of the BSE Hamiltonian, thus allowing us to still apply the formalism. Therefore, we define a small enough energy difference $\eta>0$, which spans the degenerate subspace around excitation energy $\Omega_{l}$, and get the expansion coefficients $c_{m}^{l}$, dependent on the polarization $\boldsymbol{E}$ of the pump pulse, via the optical transition matrix elements:
$c_{m}^{l}\propto\langle 0|\boldsymbol{E}\cdot\hat{\boldsymbol{v}}|m\rangle\;\delta\left(|\Omega^{(m)}-\Omega_{l}|<\eta\right).$ (18)
Here $|0\rangle$ denotes the ground state and $\hat{\boldsymbol{v}}=\mathrm{i}[\hat{H},\hat{\boldsymbol{r}}$] the single-particle velocity operator.
As a consequence, for photoemission intensity from the superposition (eigen)state $|S_{l}\rangle$, also the transition density matrix elements are linearly combined and thus Eq. II.1 must be modified accordingly ($\boldsymbol{Q}=0$):
@@
$\cdot\delta\left(\omega+\Omega_{l}-\varepsilon_{v,\boldsymbol{q}}-E_{\mathrm{kin}}\right)$ (19)
Note, that the new transition density matrix $\sum_{m}c^{l}_{m}X_{v,c,\boldsymbol{q}}^{(m)}$ can readily be identified as the BSE eigenvector of the superposition state. The procedure presented in this section can thus also be understood as a simple basis change within the subspace of degenerate eigenstates, or, in other words, as a change from the basis obtained through solution of the BSE to a basis determined by the pump polarization.
In 2D systems excitons typically occur as degenerate pairs, and only their superposition should be considered as the physical exciton, which then, for instance, also respects the symmetries of the system [41, 42]. In contrast to those eigenstates, the pump pulse in general breaks the system’s symmetry and may create superimposed exciton states that depend on the pump polarization. As we will show in Sec. III.3, such pump polarization effects have a considerable imprint on exciton photoemission.
@@ III.1 Calculation of hBN quasiparticle and optical properties
In the following, we exemplify our formalism by the simulation of excitonic photoemission in monolayer hexagonal boron nitride (hBN) as a prototypical 2D system. As a starting point, we obtain wave functions and energy levels from DFT calculations with Quantum Espresso [43, 44, 45], using the PBE-GGA exchange-correlation functional [46] and a lattice constant of 2.507 Å, see Fig. 1 (a). To compensate for the systematic underestimation of the band gap within semi-local DFT, we apply the $G_{0}W_{0}$ method to correct the energy levels. Utilizing the BerkeleyGW code [1, 2, 47], we obtain a direct (indirect) band gap of 7.6 eV (7.4 eV), which is comparable to literature values from similar methods, see e.g. Ref. [48] for a comparison. We show the $G_{0}W_{0}$ band structure in Fig. 1 (b) and detail the computational methods in Appendix C.
Figure 2: Optical excitation of monolayer hBN. (a) Independent-particle spectrum, based on a $G_{0}W_{0}$ calculation in dark blue and the BSE spectrum in pink. In the latter, $S_{1}$ and $S_{2}$ energies are marked by arrows. Panels (b) and (c) show the $G_{0}W_{0}$ band structure with overlays indicating the BSE eigenvector composition of the $S_{1}$ (blue) and $S_{2}$ (green) excitations of the system, respectively.
Subsequently, we use the DFT wave functions and $G_{0}W_{0}$ energy levels, and solve the BSE within the TDA and for $\boldsymbol{Q}=0$. This yields the excitation energies $\Omega_{m}$ and the BSE eigenvectors $|m,\boldsymbol{Q}=0\rangle\equiv|m\rangle$, which we use to simulate optical spectra via Eq. II.1. As shown in Fig. 2 (a), the BSE spectrum features a range of excitonic peaks below the band gap. These are often interpreted as hydrogen-like solutions of the Wannier equation [49], though they have also been found to deviate from that model [41]. For comparison, we include the optical spectrum computed within the independent-particle approximation, shown as a dashed blue curve in Fig. 2 (a). In the following, we focus on the two lowest excitons, located at 5.3 eV ($S_{1}$) and 6.2 eV ($S_{2}$).
Both, $S_{1}$ and $S_{2}$ are composed of two degenerate eigenstates of the BSE Hamiltonian. This results in an additional dependence on the pump polarization, as discussed in Sec. II.2, and which will become relevant at a later stage in Sec. III.3. For $S_{1}$ at 5.3 eV, the two degenerate valence-conduction transitions occur at $K$ and $K^{\prime}$, respectively, with their momentum-dependent transition weights indicated in Fig. 2 (b). The situation is similar for $S_{2}$ at 6.2 eV, albeit with transition weights being more spread out along the Brillouin-zone boundary, as can be seen in Fig. 2 (c).
@@ III.2 Exciton photoemission in hBN
With the quasiparticle and optical properties at hand, we now illustrate consequences of our theory for exciton photoemission from monolayer hBN. In the following, we use a 26.5 eV probe energy with a Gaussian energy broadening of 0.1 eV. For a better insight into the pump-induced effects on the momentum maps, we ignore the trivial effect of probe polarization, stemming from the $|\boldsymbol{A}\boldsymbol{k}|^{2}$ prefactor in Eq. II.1. Note, that our formulation only considers excitonic photoemission, thereby neglecting photoemission from the initially populated electronic structure and associated effects like its depopulation due to exciton creation. This decoupling is a direct consequence of energy conservation in excitonic photoemission Eq. II.1. The exciton energy $\Omega^{(m,\mathbf{Q})}$ is transferred to the photoelectron, yielding higher kinetic energies than those obtained from photoemission out of the initially populated electronic states (see Fig. 3).
Figure 3: ARPES bandmap for the $S_{1}$ exciton in $k_{x}$ direction (cut of the BZ along $\Gamma-K^{\prime}-M_{2}$, see inset with three neighboring BZs). The $G_{0}W_{0}$ bandstructure is overlaid in full blue lines, with a duplicate of the valence band, shifted by $\Omega$, as a blue dashed line.
In Fig. 3 we show the simulated bandmap for photoemission from the $S_{1}$ exciton together with the $G_{0}W_{0}$ bandstructure (full blue lines) along the $\Gamma-K^{\prime}-M_{2}$ path (corresponding to the $k_{x}$ direction, see inset). According to Eq. II.1, the energy dispersion of exciton photoemission is determined by the hole component of the exciton. Consequently, the exciton’s trARPES signature appears at the photoelectron kinetic energy corresponding to the valence band energy plus the pump energy $\Omega$ (blue dashed line). This finding, which is well-established in the literature [50, 19], enables the estimation of the exciton’s spatial extent from photoemission data under the assumption of an $s$-like envelope function [4, 5, 51].
@@
Next, it is expedient to compare momentum maps of exciton photoemission with those for photoemission from the ground state, which we obtain by the standard way of simulating POT [52]. In Fig. 4 (a), we present such a map for a kinetic energy slightly below the valence band maximum (VBM) ($E-E_{\mathrm{VBM}}=-0.2$ eV), which is comprised of six triangular features, each around the $KK^{\prime}$ points. This pattern is analogous to the case of graphene, where the triangular features are also broken in the outgoing direction of the Brillouin-zone, leading to croissant- or horseshoe-like shapes. This “dark corridor” is due to matrix element effects in the photoemission process and has already been described in great detail for the case of graphene [53, 54, 14]. When assuming a static, charge-transfer induced population of the conduction band, the latter would also be accessible to photoemission from the ground state, which we simulate in panel (b) of Fig. 4, using a kinetic energy of 26.6 eV, which is above the conduction band minimum (CBM) ($E-E_{\mathrm{CBM}}=+0.6$ eV). In contrast to the valence band case depicted in panel (a), the features around the Dirac point are now elongated in the directions going to the $MM^{\prime}$ points, which is reasonable when looking at the bandstructure (see e.g. Fig. 1). Remarkably, the dark corridor from the matrix element effect is now pointing to the inside of the BZ. Returning now to the simulation of exciton photoemission, we show a momentum map of the $S_{1}$ exciton in Fig. 4 (c), using a kinetic energy of 23.9 eV and a pump polarization as indicated by a black arrow. As already discussed in the preceding paragraph, the exciton inherits the energy dispersion of the hole (here: the highest valence band), but, in light of Eq. II.1, follows the momentum distribution given by the Fourier-transformed electron part of the exciton wave function (here: the lowest conduction band). As a direct imprint of the latter, the dark corridor is also pointing to the inside of the BZ, which we can see clearly in Fig. 4 (c). Having established the unoccupied part of the electronic structure as another direct marker of exciton photoemission, we will now examine the different contributions to the momentum maps in more detail.
@@
Figure 5: Decomposition of exciton ARPES maps for $S_{1}$ in row (a) and $S_{2}$ in row (b), with black arrows indicating the polarization of the pump laser. First column: Fourier-transformed conduction-band wave functions. Second column: energy-conservation delta-function. Third column: BSE eigenvectors. Fourth column: resulting exciton ARPES maps as products of first to third columns. Each map is individually normalized to one. For $S_{1}$ in row (a), the horseshoes point away from $\Gamma$, with minor modulations along their contours. For $S_{2}$ in row (b), four of the six horseshoes are rotated, when compared to the first exciton. This is a direct result of the differences in the momentum-space distribution of their BSE eigenvectors.
@@
Using Eq. II.1, we dissect the photoemission maps for both excitons, $S_{1}$ and $S_{2}$, in order to understand the influence of each contributing factor. Since—in this case—both excitons only involve transitions from the highest valence to the lowest conduction band, the sums over $v$ and $c$ in Eq. II.1 collapse. Consequently, we are left with a single product of $|\mathcal{F}[\xi]|^{2}$, $\delta_{E}$ and $|X|^{2}$, involving only wave functions from the lowest conduction band, energies of the highest valence band and BSE transition coefficients from the highest valence band to the lowest conduction band. Fig. 5 shows these three factors, as well as the full photoemission maps, for the two excitons in row (a) ($S_{1}$) and row (b) ($S_{2}$), respectively. The direction of the pump polarization is indicated in the rightmost column. For both excitons $S_{1}$ and $S_{2}$, the triangular patterns are generated by the $\delta_{E}$ factor, with $|\mathcal{F}[\xi]|^{2}$ only leading to slight modifications. The maps in the first two columns, $|\mathcal{F}[\xi]|^{2}$ and $\delta_{E}$, are very similar between the two excitons $S_{1}$ and $S_{2}$, with minor differences arising due to the unequal photoelectron kinetic energies (23.9 eV for $S_{1}$ and 24.8 eV for $S_{2}$). Since all other areas already have vanishing intensity, the additional multiplication with the BSE eigenvector can at most lead to intensity differences within the area of the horseshoes. As can be seen from the third column, the BSE eigenvectors of the two excitons differ, leading to the differences in their respective photoemission maps (rightmost column of Fig. 5). In the following, those differences are inspected more closely.
@@
For $S_{1}$, the BSE eigenvector (third column in Fig. 5 (a)) mostly exhibits a six-fold symmetry. It is centered around the $K$ and $K^{\prime}$ points and varies weakly along the contours of the horseshoes, thus leading only to small intensity variations (fourth column in Fig. 5 (a)) along them. Contrary, in the case of $S_{2}$ in row (b), the BSE eigenvector exhibits a nodal line perpendicular to the pump polarization — which has broken the six-fold symmetry — while being non-zero along the remaining four sides of the hexagonal BZ. Since the pump polarization was chosen along a high-symmetry direction (indicated by black arrows in the rightmost column), it acts as a symmetry axis for the BSE eigenvector and, consequently, for the resulting photoemission map. In contrast to the case of $S_{1}$, the regions around $K$ and $K^{\prime}$ show much more details for $S_{2}$, particularly featuring small nodal lines, which, upon multiplication of the BSE eigenvector with the horseshoe pattern produced by $|\mathcal{F}[\xi]|^{2}$ and $\delta_{E}$, suppress photoemission intensity. Moreover, the local maxima are shifted off $K$ and $K^{\prime}$. As can be seen in the last column of Fig. 5, the combination of all three factors effectively leads to a redistribution of spectral intensity within the horseshoes. By selectively accessing the different factors in Eq. II.1, we are thus able to deconvolute different contributions to the momentum maps. In the general case of multiple transitions within an exciton, however, the sums over different valence and conduction bands remain and, as a result, the above analysis would become less straightforward, especially due to the coherent summation of the electron part. Since in the present case of the hBN example we found the most asymmetry to arise due to the BSE eigenvectors in conjunction with the pump polarization, we will discuss this effect in more detail in the following.